OpenStudy (anonymous):

integral help! (cal 2) posted below! thanks in advance!

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ 3\sec ^{2}t }{ 6+3\tan t }dt\]

OpenStudy (anonymous):

put tan t=x sec^2 t dt=dx

OpenStudy (anonymous):

I think my first step should be a substitution. so im thinking the bottom somehow

OpenStudy (anonymous):

ahh ok let me start with this

OpenStudy (anonymous):

first take common 3 from numerator and denominator and cancel ,then substitute as above.

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dx }{ 2+x }\]

OpenStudy (anonymous):

correct now integrate.

OpenStudy (anonymous):

i take another substitution?

OpenStudy (anonymous):

no need when numerator is diff. of denominator \[\int\limits \frac{ f'(x) }{ f(x) }dx=\ln \left| f(x) \right|\]

OpenStudy (anonymous):

sorry, i am just learning this and these are some of the class examples i have written down

OpenStudy (anonymous):

ooooh! makes sense because the derivative of lnx is 1/x